Save Question 14 (1 point) If a motor has three times as much power as another, then the less power motor A) can do the same work in three times the time O B) can never do the same work as the larger motor O c) can do the same work in the same time C D) can do the same work in one-third the time Save Question 15 (1 point)
Answer #1

let power of ist motor ,P1 = 3 P

t1 = time of ist motor = t

power of 2nd motor ,P2 = P

t2 = time of 2nd motor = 3t

--------------------

power = work / time

W1 (work done by ist motor) = P1*t1

W1 = 3P*t ----------------------1

W2(work done by 2nd motor) = P2*t2

W2 = P*3t = 3Pt --------------------2

from eq 1 and 2 less power motor can do the same work in three times the time.